OpenStudy (anonymous):

satelitte i cant seem to acces the convo we were having..i tried fidning the x interecpt by making it equal zero but im getting no solutions:S...n what did u say about the points of inflection r they none?

OpenStudy (anonymous):

\[x-\frac{256}{x^2}=0\] \[\frac{x^3-256}{x^2}=0\] \[x^3-256=0\] \[x^3=256\] \[x=\sqrt[3]{256}\]

OpenStudy (anonymous):

so the point would be about (85,0)